# Eugene Kritchevski

### v1

@@ 3 Reduction to Two Dimensions in the General Case
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
The key point is that replacing $(x_{i},y_{i})$ by $(x_{i},y_{i}^{*})$ does not decrease $F_{d}$, does not change $s^{\pm}(x,y)$, and forces the ratio $y_{i}^{*}/x_{i}$ to be extremal. The frozen extremal indices in $I_{-}\cup I_{+}$ ensure that the original extremal ratios are preserved throughout.
We now apply this replacement successively to all $i\in I_{0}$, always leaving previously modified coordinates unchanged. This yields a vector $y^{*}$ such that
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),$
and every ratio $y_{i}^{*}/x_{i}$ is extremal:
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
With $y^{*}$ fixed, we now apply exactly the same coordinatewise extremal-pushing procedure to the pair $(u,v)$. We freeze the indices where the ratios $v_{i}/u_{i}$ already attain the extremal values $b^{\pm}$, and modify only the remaining coordinates exactly as above: for each active coordinate we replace $v_{i}$ by either $b^{-}u_{i}$ or $b^{+}u_{i}$, according to the monotonicity of $F_{d}(x,y^{*};u,v)$ in $v_{i}$. This produces a vector $v^{*}$ such that
$s^{\pm}(u,v^{*})=s^{\pm}(u,v),\qquad F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y^{*};u,v),$
and each ratio $v_{i}^{*}/u_{i}$ is extremal:
$\frac{v_{i}^{*}}{u_{i}}\in\{\,b^{-},\,b^{+}\,\}.$
This completes the proof. ∎
Lemma 3.2 ($2$-point support collapse).
Assume that the coordinate ratios of $y^{*}/x$ take only the two extremal values $a^{\pm}$, and the coordinate ratios of $v^{*}/u$ take only the two extremal values $b^{\pm}$, where
$a^{\pm}=s^{\pm}(x,y^{*}),\qquad b^{\pm}=s^{\pm}(u,v^{*}).$
Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;F_{d}(x,y^{*};u,v^{*}),$
and
$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}.$
Proof.
If $a^{+}=a^{-}=a$, then $y^{*}=ax$ and hence $F_{d}(x,y^{*};u,v^{*})=1$. The lemma then holds, for example, with
$x^{\prime}=(1,1),\qquad y^{\prime}=(a,a),\qquad u^{\prime}=(1,1),\qquad v^{\prime}=(b^{-},b^{+}).$
We now treat the nontrivial case $a^{+}>a^{-}$. Define
$I_{+}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{+}\bigr\},\qquad I_{-}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{-}\bigr\},$
which form a partition of $\{1,\dots,d\}$. Set
$x_{\pm}:=\sum_{i\in I_{\pm}}x_{i},\qquad u_{\pm}:=\frac{1}{x_{\pm}}\sum_{i\in I_{\pm}}x_{i}u_{i},$
so that
$x_{\pm}u_{\pm}=\sum_{i\in I_{\pm}}x_{i}u_{i}.$
Set
$y_{\pm}:=a^{\pm}x_{\pm},\qquad v_{\pm}:=b^{\pm}u_{\pm},$
and define the two-dimensional vectors
$x^{\prime}:=(x_{-},x_{+}),\qquad y^{\prime}:=(y_{-},y_{+}),\qquad u^{\prime}:=(u_{-},u_{+}),\qquad v^{\prime}:=(v_{-},v_{+}).$
A direct computation using the above definitions gives
$x^{\prime}\cdot u^{\prime}=x_{-}u_{-}+x_{+}u_{+}=\sum_{i\in I_{-}}x_{i}u_{i}+\sum_{i\in I_{+}}x_{i}u_{i}=x\cdot u,$
and similarly
$x^{\prime}\cdot v^{\prime}=x\cdot v^{*},\qquad y^{\prime}\cdot u^{\prime}=y^{*}\cdot u,\qquad y^{\prime}\cdot v^{\prime}=y^{*}\cdot v^{*}.$
Hence
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=\frac{(x^{\prime}\cdot u^{\prime})(y^{\prime}\cdot v^{\prime})}{(x^{\prime}\cdot v^{\prime})(y^{\prime}\cdot u^{\prime})}=\frac{(x\cdot u)(y^{*}\cdot v^{*})}{(x\cdot v^{*})(y^{*}\cdot u)}=F_{d}(x,y^{*};u,v^{*}),$
which completes the proof.
@@
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
@@
Apply lemma 3.1 to obtain $y^{*},v^{*}$ with
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
then lemma 3.2 to $(x,y^{*};u,v^{*})$ to obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=F_{d}(x,y^{*};u,v^{*}),$
which gives the claim. ∎
@@
By definition, $R(AB)$ is the maximum distortion among all $2\times 2$ sub-blocks of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ such that this maximum is attained.
Let
@@
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then $R(A^{\prime})=\operatorname{Dist}(x,y)=\alpha$ and $R(B^{\prime})=\operatorname{Dist}(u,v)=\beta$. Then
$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix}$,
and therefore
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;R(A^{\prime}B^{\prime}).$
Combining the above steps and using that $\Phi$ is increasing in each variable gives
@@
$\leq\Phi(\alpha,\beta)$
$\leq\Phi(R(A),R(B)).$
This completes the proof.
@@ 4 Distortion of Products
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and define the partial products
$P_{n}:=A_{n}\cdots A_{1}.$
The matrices may have arbitrary dimensions, provided the products are well defined.
@@
Since $\Psi$ is increasing in each variable, distortion bounds propagate along the product. If
@@
With $p$ fixed, this iterates a single Möbius map. Writing
@@
its $n$th iterate can be expressed as
@@
For convenience, we record the following standalone theorem.
@@ 5 Comparison with the Birkhoff–Bushell
$d_{H}(x,y)=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
@@
In this log setting, our envelope inequality becomes
@@
$\Theta(h)=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p=\sqrt{R(A)}.$
In the nontrivial case $p>1$, a direct computation shows that $\Theta$ is strictly increasing, strictly concave, and satisfies
@@
Thus $\Theta$ lies below its tangent line at the origin:
@@
which yields the classical BB contraction:
@@
so the envelope bound $\Theta(h)$ remains uniformly bounded, whereas the BB linear bound
$\kappa(A)\,h$
diverges. Thus both bounds agree to first order at $h=0$ (and therefore yield the same Lyapunov–gap asymptotics for distortion of high powers of $A$), but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp in the non-asymptotic regime.
For illustration, Figure 2 compares the nonlinear envelope map $h\mapsto\Theta(h)$ with the classical BB line $k(A)h$, together with the saturation level $\log R(A)$.
@@
Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$. The envelope contraction saturates at $\log R(A)$, while the BB line grows without bound, showing that the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$ for large $h$.
Historical perspective.
Birkhoff’s fundamental insight was that positive operators act as contractions in the Hilbert metric, allowing the Frobenius–Perron alignment phenomenon to be deduced from the Banach fixed-point theorem. Bushell later refined this viewpoint by computing an explicit contraction constant. In most contraction theorems such constants are technically useful but algebraically unremarkable: they serve as quantitative bounds and carry little intrinsic meaning. Bushell’s constant is a striking exception. Its closed form,
$\kappa(A)=\frac{\sqrt{R(A)}-1}{\sqrt{R(A)}+1},$
is unexpectedly symmetric and rigid—far too elegant, in retrospect, to be merely a sharp Lipschitz coefficient. Moreover, the metric framework of the Hilbert metric plays no essential role in the classical proof: it is used only to invoke the Banach fixed-point theorem, not to explain the mechanism of alignment itself. This indicates that the contraction phenomenon behind the Frobenius–Perron theory is not inherently metric but fundamentally algebraic, encoded in the behaviour of ratios rather than in the geometry induced by the Hilbert metric.
This motivates our approach: the envelope inequality shows that worst-case distortion evolves under Möbius iterations with a common attracting fixed point at $1$. In this formulation, alignment arises from the explicit fixed-point dynamics of these maps rather than from a geometric contraction in the Hilbert metric.
6 Discussion and outlook
The envelope inequality is dimension-free, and the same collapse mechanism extends to a broad class of infinite matrices and positive operators on cones. The operator-theoretic generalizations will be developed elsewhere.
A second direction concerns other projective quantities. Since the collapse argument in Section 3 uses only the four-point ratios
$\frac{(x\!\cdot\!u)(y\!\cdot\!v)}{(x\!\cdot\!v)(y\!\cdot\!u)},$
it is natural to conjecture that analogous two-dimensional collapse mechanisms may apply to a broader class of projective functionals.
The envelope inequality also suggests applications in Markov chain theory. Classical asymptotic mixing bounds typically rely on Doeblin, Dobrushin, or Hilbert-metric methods, whereas the envelope provides an exact non-asymptotic contraction profile. This may offer an alternative route to quantitative bounds for finite-time mixing of positive stochastic matrices.
The recursion in Section 4 provides a deterministic model for worst-case distortion in products of positive matrices; its transfer-matrix form makes the evolution completely explicit. It may also serve as a foundation for combining our methods with probabilistic techniques to study non-asymptotic behavior of random products.
These directions lie beyond the scope of the present paper. Our purpose here is purely foundational: to isolate and prove the core envelope inequality in its simplest form.
Appendix A Calculus Tools
In this appendix we prove properties of the functions (for $\alpha,\beta\geq 1$)
$f_{\alpha,\beta}(t)=\frac{(1+t)(1+\alpha\beta\,t)}{(1+\alpha t)(1+\beta t)},\qquad t>0,$
and
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
Lemma A.1.
The function $f_{\alpha,\beta}$ has the following properties:
  1. 1.
    $f_{\alpha,\beta}(t)\geq 1$ for all $t>0$;
  2. 2.
    if $\alpha,\beta>1$, the unique global maximum is attained at
    $t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
    and
    $f_{\alpha,\beta}(t)\leq f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta);$
  3. 3.
    if $\alpha=1$ or $\beta=1$, then $f_{\alpha,\beta}(t)\equiv 1$.
Proof.
(1) Expanding $f_{\alpha,\beta}(t)\geq 1$ gives
$(1+t)(1+\alpha\beta t)\;\geq\;(1+\alpha t)(1+\beta t),$
which simplifies to
$(\alpha-1)(\beta-1)t\geq 0,$
valid for all $\alpha,\beta\geq 1$ and $t>0$.
(2) Set $\varphi(t)=\log f_{\alpha,\beta}(t)$. A straightforward algebraic simplification yields
$\varphi^{\prime}(t)=\,\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
Thus $\varphi^{\prime}(t)=0$ exactly at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
and this is the unique global maximum. Substituting $t^{*}$ gives the stated value of $f_{\alpha,\beta}(t^{*})$.
(3) If $\alpha=1$ or $\beta=1$, the formula reduces to $f_{\alpha,\beta}(t)\equiv 1$. ∎
Lemma A.2.
For $\alpha,\beta\geq 1$, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable: for all $\alpha_{1}\leq\alpha_{2}$ and $\beta_{1}\leq\beta_{2}$,
$\Phi(\alpha_{1},\beta)\leq\Phi(\alpha_{2},\beta),\qquad\Phi(\alpha,\beta_{1})\leq\Phi(\alpha,\beta_{2}).$
Proof.
Write $p=\sqrt{\alpha}$ and $q=\sqrt{\beta}$, so $p,q\geq 1$, and define
$\psi(p,q)=\frac{1+pq}{p+q},$
so that $\Phi(\alpha,\beta)=\psi(p,q)^{2}$. For fixed $q\geq 1$ we compute
$\frac{\partial\psi}{\partial p}=\frac{q(p+q)-(1+pq)}{(p+q)^{2}}=\frac{q^{2}-1}{(p+q)^{2}}\;\geq\;0.$
Thus $\psi$ is increasing in $p$, and therefore so is $\Phi=\psi^{2}$. By symmetry the same argument applies to $q$. ∎

### v2

@@ 3 Reduction to Two Dimensions
[…]
Since $\tilde{b}^{\pm}\in[b^{-},b^{+}]$ and $f$ is decreasing in its first argument and increasing in its second, we conclude that
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=f(b^{-},b^{+})\;\geq\;f(\tilde{b}^{-},\tilde{b}^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},\tilde{v})=F_{d}(x,y;u,v).$
This completes the proof.
@@
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
@@
Apply Lemma 3.1 to obtain $y^{*},v^{*}$ with $F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y;u,v)$. Then apply Lemma 3.2 to $x,y^{*},u$ and $v^{*}$ and obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y^{*};u,v^{*})$. ∎
@@
By definition, $R(AB)$ is attained by a $2\times 2$ sub-block of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ attaining this maximum. Let
@@
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then
$R(A^{\prime})=\operatorname{Dist}(x^{\prime},y^{\prime})=\alpha,\qquad R(B^{\prime})=\operatorname{Dist}(u^{\prime},v^{\prime})=\beta,$
and
$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix},\qquad\text{so}\qquad F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=R(A^{\prime}B^{\prime}).$
Using Theorem 2.2 and the monotonicity of $\Phi$,
@@
$\leq\Phi\bigl(R(A^{\prime}),R(B^{\prime})\bigr)$
$=\Phi(\alpha,\beta)$
$\leq\Phi\bigl(R(A),R(B)\bigr).$
@@ 4 Distortion of Products
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and set
$P_{n}:=A_{n}\cdots A_{1},$
assuming the products are well defined.
@@
Since $\Psi$ is increasing, distortion bounds propagate along products. If
@@
This recursion is the iteration of a single Möbius map. Writing
@@
we obtain the closed-form expression
@@
Collecting the preceding arguments, we record the following theorem.
@@ 5 Comparison with the Birkhoff–Bushell
$d_{H}(x,y):=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
@@
In this logarithmic setting, the envelope inequality takes the form
@@
$\Theta(h):=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p:=\sqrt{R(A)}.$
A direct computation gives
$\Theta^{\prime}(h)=\frac{(p^{2}-1)e^{h/2}}{(1+pe^{h/2})(p+e^{h/2})},\qquad\Theta^{\prime\prime}(h)=-\frac{(p^{2}-1)e^{h/2}p(e^{h}-1)}{2(1+pe^{h/2})^{2}(p+e^{h/2})^{2}}.$
In the nontrivial case $p>1$, it follows that $\Theta$ is increasing and concave on $0,\infty)$, and satisfies
@@
By concavity, $\Theta$ lies below its tangent line at the origin,
@@
which yields the classical BB contraction estimate
@@
so the envelope bound remains uniformly bounded, whereas the linear BB bound $\kappa(A)h$ diverges. The two bounds therefore agree to first order at $h=0$, but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp for all $h$.
Figure 2 compares the nonlinear envelope $h\mapsto\Theta(h)$ with the classical BB line $\kappa(A)h$ and the saturation level $\log R(A)$.
@@
Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$. The envelope saturates at $\log R(A)$, while the BB bound grows without bound.